OpenStudy (sodapop):

The population of a local species of mosquitos can be found using an infinite geometric series where a1 = 740 and the common ratio is one sixth. Write the sum in sigma notation and calculate the sum (if possible) that will be the upper limit of this population.

OpenStudy (sodapop):

@Hero

OpenStudy (sodapop):

the summation of 740 times one sixth to the i minus 1 power, from i equals 1 to infinity. ; the sum is divergent the summation of 740 times one sixth to the i minus 1 power, from i equals 1 to infinity. ; the sum is 888 the summation of 740 times one sixth to the i power, from i equals 1 to infinity. ; the series is divergent the summation of 740 times one sixth to the i power, from i equals 1 to infinity. ; the sum is 88

OpenStudy (sodapop):

@SolomonZelman

OpenStudy (phi):

any idea what this means a1 = 740 ?

OpenStudy (sodapop):

Im guessing it goes in some kind of formula

OpenStudy (phi):

a1 is short hand for the name of the first term. an infinite geometric series is \[ a_1 + a_2+ a_3 + ...\] which goes on forever. they are telling you the first term is 740

OpenStudy (phi):

the common ratio is one sixth that means the ratio of two consecutive terms is 1/6. For example \[ \frac{a_2}{a_1} = \frac{1}{6} \] if we "solve" for a2 we get \[ a_2 = \frac{1}{6} \cdot 740 \]

OpenStudy (phi):

also, (because the ratio of two consecutive terms is 1/6 ) \[ \frac{a_3}{a_2} = \frac{1}{6} \] if we solve for a3 we get \[ a_3 = \frac{1}{6}\cdot a_2 = \frac{1}{6}\cdot \frac{1}{6} a_1 = 740 \left(\frac{1}{6}\right)^2 \]

OpenStudy (phi):

if we use those numbers in \[ a_1 + a_2 + a_3 + ...\] we get \[ 740 + 740 \cdot \frac{1}{6} + 740 \cdot \left(\frac{1}{6} \right)^2+... \]

OpenStudy (sodapop):

Alright

OpenStudy (sodapop):

So it would look like this \[\sum_{i=1}^{\infty}=740(\frac{ 1 }{ 6 })^{i}\]

OpenStudy (sodapop):

@phi

OpenStudy (phi):

the sigma (Greek capital S, short for sum) you posted has i=1 at the bottom that means the first term has i=1. in other words, \[ 740 \left(\frac{1}{6}\right)^1 \] but you want \[ 740 + 740 \cdot \frac{1}{6} + 740 \cdot \left(\frac{1}{6} \right)^2+... \\ 740 \cdot \left(\frac{1}{6} \right)^0+ 740 \cdot \left(\frac{1}{6} \right)^1 + 740 \cdot \left(\frac{1}{6} \right)^2+... \] notice that (1/6) to the zero power is 1. anything to the 0 exponent is 1 does that pattern help you figure out what the sigma notation looks like ?

OpenStudy (phi):

See this for if this series converges: https://en.wikipedia.org/wiki/Geometric_series#Common_ratio

OpenStudy (sodapop):

\[\sum_{i=1}^{\infty}(\frac{ 1 }{ 6 })^{i=1}\]'the sum is divergent \[\sum_{i=1}^{\infty}(\frac{ 1 }{ 6 })^{i=1}\];the sum is 88 \[\sum_{i=1}^{\infty}(\frac{ 1 }{ 6 })^{i}\];the series is divergent \[\sum_{i=1}^{\infty}(\frac{ 1 }{ 6 })^{i=1} \];the sum is 88

OpenStudy (sodapop):

These are my answers @phi

OpenStudy (sodapop):

Which would be the correct one ? @Hero

OpenStudy (sodapop):

@Jim766